You are back on the HOA board. The time has come to buy more trees and bushes for the neighborhood park. The first order was for 24 bushes and 16 trees and the bill totaled $3960. The second order was 12 bushes and 32 trees and the bill totaled $5220.
Write an equation illustrating the first bill using b for the cost of the bushes and t for the cost of the trees
Write an equation illustrating the second bill using b for the cost of the bushes and t for the cost of the trees
Using the two equations from above and elimination method to find the cost of each bush and the cost of each tree.

Answer by stanbon(75887)

You are back on the HOA board. The time has come to buy more trees and bushes for the neighborhood park.
The first order was for 24 bushes and 16 trees and the bill totaled $3960.
24b + 16t = 3960
---
The second order was 12 bushes and 32 trees and the bill totaled $5220.
12b + 32t = 5220
-------
Write an equation illustrating the first bill using b for the cost of the bushes and t for the cost of the trees
Write an equation illustrating the second bill using b for the cost of the bushes and t for the cost of the trees
Using the two equations from above and elimination method to find the cost of each bush and the cost of each tree.
---
Multiply thru the 2nd eq. by 2:
24b + 16t = 3960
24b + 64t = 10440
---
Subtract and solve for "t":
48t = 6480
t = 135 (# of trees purchased)
---
solve for "b":
24b + 16*135 = 3960
24b = 1800
b = 75 (# of bushes purchased)
==================================
Cheers,
Stan H.